An ideal transformer with 300 turns in its primary and 500 turns in its secondary has a peak primary voltage of 12 volts. Its peak secondary voltage is

a. 10

b. 20

c. 30

A transformer's primary and secondary coils are connected so that the current can flow directly from one to the other.

a. true

b. false
Answer #1

1)
Ns/Np = Vs/Vp
Ns - Number of turns in the secondary coil
Np - Number of turns in the primary coil
Vs - Voltage in the secondary coil
Vp - Voltage in the primary coil

Vs = Vp * (Ns/Np)
= 12 * (500/300)
= 20 V

2)
False
Current does not flow directly, the current is induced indirectly.